Alex bought 10 potatoes and 6 onions for $15. Carlos bought 5 potatoes and 2 onions for $6. What is the cost of the onions?


1
Expert's answer
2021-06-07T17:17:01-0400

Let "x=" the cost of the onion, "y=" the cost of the potatoe. Then



"10x+6y=15"

"5x+2y=6"


"10x+6y-2(5x+2y)=15-2(6)"

"2y=3"

"y=1.50"

The cost of the onions is $1.50 per unit.